# Off-policy Evaluation with Deeply-abstracted States

### v2

@@ 4.2 Model-based irrelevance conditions
Forward MDP-based model-irrelevance. We first discuss the connections between the model-irrelevance given in Definition 3 and the notions of $Q^{\pi}$-, $\rho^{\pi}$- and $w^{\pi}$-irrelevance, and introduce the following conditions to establish Fisher consistency. These conditions are mild and frequently imposed in the RL and OPE literature (see e.g., Thomas and Brunskill,, 2016; Kallus and Zhou,, 2018; Chen and Jiang,, 2019; Fan et al.,, 2020; Cai et al.,, 2021; Shi et al.,, 2021; Kallus and Uehara,, 2022).
Assumption 1 (Boundedness)
All immediate rewards are uniformly bounded.
Assumption 2 (Coverage)
The denominator in (2) is strictly positive.
Assumption 3 (Stationary)
The MDP $(S_{t},A_{t},R_{t})_{t\geq 1}$ is stationary over time.
The following lemma summarizes the findings. Results in the first two bullet points are based on those in the existing literature (see e.g., Li et al.,, 2006; Pavse and Hanna,, 2023).
Lemma 1 (OPE under model-irrelevance)
Assume Assumptions 13 hold. Let $\phi$ denote a model-irrelevant abstraction. Suppose $\phi$ is additionally $\pi$-irrelevant. Then:
  • Fisher consistency of SIS: While $\phi$ is not necessarily $\rho^{\pi}$-irrelevant, SIS (Method 2) with a history-dependent IS ratio (detailed in the proof of Lemma 1 in Appendix D.1) is Fisher consistent when applied to the abstract space;
The last three bullet points validate the SIS, MIS and DRL estimators, despite $\phi$ being neither $w^{\pi}$-irrelevant nor $\rho^{\pi}$-irrelevant. By definition, $\rho^{\pi}$-irrelevance can be achieved by selecting state features that adequately predict the IS ratio. However, methods for constructing $w^{\pi}$-irrelevant abstractions remain less clear. In the following, we introduce a backward MDP model-based irrelevance condition that ensures $w^{\pi}$-irrelevance.
Backward MDP-based model-irrelevance. To illustrate the rationale behind the proposed model-based abstraction, we introduce the backward MDP model by reversing the time index. Under the (forward) MDP model assumption described in Section 3.1 and that the behavior policy $b$ is not history-dependent, actions and states following $S_{t}$ are independent of those occurred prior to the realization of $S_{t}$. Accordingly, $(S_{t-1},A_{t-1})$ is conditionally independent of $\{(S_{k},A_{k})\}_{k>t}$ given $S_{t}$. Recall that $T$ corresponds to the termination time of trajectories in $\mathcal{D}$. We define a time-reversed process consisting of state-action-reward triplets $\{(S_{t},A_{t},b(A_{t}|S_{t})):t=T,\dots,1\}$. Its dynamics is described as follows (see also Figure 2(b) for the configuration):
@@
Given this MDP, analogous to Definition 3, our objective is to identify a state abstraction that is crucial for predicting the reward (e.g., the behavior policy) and the reversed transition function. We provide the formal definition of the proposed backward MDP-based model-irrelevance (short for backward-model-irrelevance) below.
@@
The conditions of backward-model-irrelevance are similar to those specified for model-irrelevance outlined in Definition 3. Equation (4) essentially requires behavior-policy-irrelevance, or reward-irrelevance in the backward MDP. Equation (5) corresponds to the “backward-transition-irrelevance”, and is equivalent to the conditional independence assumption between the pair $(A_{t},\phi(S_{t}))$ and $S_{t+1}$ given $\phi(S_{t+1})$. As previously assumed, $S_{t}$ can be decomposed into the union of relevant features $\phi(S_{t})$ and irrelevant features $\psi(S_{t})$ (see Figure 1), leading to the following factorization:
$\mathbb{P}(S_{t+1}=s^{\prime}|A_{t},\phi(S_{t}))=\mathbb{P}(\psi(S_{t+1})=\psi(s^{\prime})|\phi(S_{t+1})=\phi(s^{\prime}))\mathbb{P}(\phi(S_{t+1})=\phi(s^{\prime})|A_{t},\phi(S_{t})).$
This indicates a two-step transition in the forward model: initially from $(\phi(S_{t}),A_{t})$ to $\phi(S_{t+1})$, and then from $\phi(S_{t+1})$ to $\psi(S_{t+1})$. Importantly, the generation of $\psi(S_{t+1})$ in the second step is conditionally independent of $A_{t}$ and $\phi(S_{t})$. Consequently, $\phi$ extracts state representations that are influenced either by past actions or past relevant features; see Figure 1(b) for an illustration. Combined with $\pi$-irrelevance and behavior-policy-irrelevance, this ensures that all information contained within the historical IS ratios $\{\rho^{\pi}(A_{k},S_{k})\}_{k<t}$ can be effectively summarized using a single $A_{t-1}$ and the abstract state $\phi(S_{t-1})$, thus achieving $w^{\pi}$-irrelevance (see Theorem 1 below).
Theorem 1 (OPE under backward-model-irrelevance)
Assume Assumptions 13 hold, and $\phi$ is both backward-model-irrelevant and $\pi$-irrelevant.
  • Then $\phi$ is both $\rho^{\pi}$-irrelevant and $w^{\pi}$-irrelevant.
@@
Theorem 1 validates the four OPE methods when applied to the abstract state space. To conclude this section, we draw a connection between the proposed backward-model-irrelevant abstraction for OPE and the Markov state abstraction (MSA) developed by Allen et al., (2021) for policy learning. MSA impose two conditions: (i) inverse-model-irrelevance, which requires $A_{t}$ to be conditionally independent of $S_{t}$ and $S_{t+1}$ given $\phi(S_{t})$ and $\phi(S_{t+1})$; (ii) density-ratio-irrelevance, which requires $\phi(S_{t})$ to be conditionally independent of $S_{t+1}$ given $\phi(S_{t+1})$. For effective policy learning, MSA requires both conditions to hold in data generating processes following a diverse range of behavior policies. When restricting them to one behavior policy, the two conditions are closely related to our backward-model-irrelevance. In particular, they imply our proposed backward-transition-irrelevance condition in (5) whereas backward-transition-irrelevance in turn yields their density-ratio-irrelevance. This allows us to adapt their algorithm to train state abstractions that satisfy our proposed backward-model-irrelevance; see Appendix B.2 for details.
Finally, it is worth noting that both the proposed backward-model-irrelevant abstraction and MSA require the behavior policy to be Markov, independent of the past observations.The following section will relax this condition and extend our proposal to accommodate history-dependent behavior policies.
Figure 3: Illustrations of the iterative procedure.
4.3 Iterative procedure for deep state abstraction
To summarize, we have reviewed the model-irrelevance condition and proposed a new backward-model-irrelevance condition. Both lead to Fisher consistent OPE estimators when confined to the corresponding abstract state spaces. This motivates us to combine the two procedures for a more condensed state abstraction, resulting in the following iterative algorithm (see Figure 3 for a visualization):
  1. 1.
    Forward abstraction: learn an abstraction $\phi_{1}$ from the ground state space $\mathcal{S}=\mathcal{X}_{0}$ to $\mathcal{X}_{1}$ using the data triplets $(S,A,R)$ that is both (forward)-model-irrelevant and $\pi$-irrelevant.
  2. 2.
    Backward abstraction: Learn an abstraction $\phi_{2}$ from the abstract state space $\mathcal{X}_{1}$ to $\mathcal{X}_{2}$ using the data triplets $(\phi_{1}(S),A,R)$ that is both backward-model-irrelevant and $\pi$-irrelevant.
  3. 3.
    Iterate the two steps to compute the final abstraction $\phi_{K}\circ\cdots\circ\phi_{2}\circ\phi_{1}$ from the ground space $\mathcal{S}$ to $\mathcal{X}_{K}$ where $K$ denotes the number of iterations and $\circ$ denotes the function composition operator.
In particular, our approach alternates between forward and backward abstraction on the state space obtained from the previous iteration. Each iteration guarantees that the cardinality of the state space does not increase, effectively maintaining or reducing complexity. Consequently, such an iterative procedure progressively reducing state cardinality, which ultimately yields a deeply-abstracted state. We thus refer to our approach as deep state abstraction (DSA).
To elaborate the usefulness of DSA in reducing state cardinality, we analyze two examples: a bandit example and an MDP example. In both examples, we focus on a specific type of state abstraction known as variable selection, which selects a sub-vector from the original state. Additionally, we focus on the class of state-agnostic target policies where $\pi$ is independent of the states. This type of policy is prevalent in causal inference and A/B testing, where the objective is to learn the global treatment effect of applying either a new or old policy consistently over time (see e.g., Hu and Wager,, 2022; Leung,, 2022; Bojinov et al.,, 2023; Shi et al.,, 2023; Xiong et al.,, 2024).
Figure 4: Two illustrative examples.
Example I (A bandit example): To illustrate the main idea, we start by considering the contextual bandit setting (CB, Dudík et al.,, 2014), which can be regarded as a special MDP with independent state transitions. Under this setting, the states are i.i.d. generated, and model-irrelevance is reduced to reward-irrelevance whereas the proposed backward-model-irrelevance simplifies to behavior-policy-irrelevance. When specialized to variable selection in CB, our proposal is reduced to the iterative confounder selection algorithm in causal inference (see e.g., Guo et al.,, 2022); see also the review of confounder selection in Appendix A. We assume the states can be divided into three independent groups, denoted by $S_{t}^{(1)}$, $S_{t}^{(2)}$ and $S_{t}^{(3)}$, respectively. Each group influences the system differently: $S_{t}^{(1)}$ affects only the reward, $S_{t}^{(2)}$ impacts both the action and the reward, and $S_{t}^{(3)}$ solely influences the action; see Figure 4(a) for an illustration. As formally proven in Lemma D.3 (see Appendix D.5):
  • The forward abstraction selects the first two groups $S_{t}^{(1)}$ and $S_{t}^{(2)}$;
  • The proposed backward abstraction selects the last two groups $S_{t}^{(2)}$ and $S_{t}^{(3)}$;
Example II (An MDP example): We next consider an MDP with three groups of states, depicted in Figure 4(b). Key observations from this example are as follows: (i) The reward depends on the state only through the first group of variables; (ii) The evolution of the state depends only on the second group. Specifically, the second group evolves first at each time step and subsequently influences the rest two groups; (iii) The behavior policy depends only on the last group; (iv) The second group is directly influenced by the previous action. Based on these observations, we show have that:
  • According to (ii) and (iv), selecting the second group achieves both transition-irrelevance (see the second equation in 3) and backward-transition-irrelevance (see (5)).
Consequently, the forward and backward abstractions select the first two and last two groups of variables, respectively. The iterative procedure again selects their intersection $S_{t}^{(2)}$ and converges in two iterations, leading to in a smaller state space. The reader is referred to Appendix D.5 for formal justifications.
In both examples, we have demonstrated the advantage of DSA in reducing state space cardinality over non-iterative procedures. However, in more general scenarios, two challenges arise from the iterative nature of DSA: (i) After forward abstraction, the behavior policy when restricted to the abstract space can be history-dependent. This would invalidate the subsequent backward abstraction for achieving $w^{\pi}$-irrelevance. (ii) After backward abstraction, the Markov assumption might be violated. This would invalidate the subsequent forward abstraction for consistent OPE.
@@ 4.3 Iterative procedure for deep state abstraction
To address both challenges, we modify the proposed backward-model-irrelevance by employing a history-dependent behavior policy $b_{t}(a_{t}|s_{t},a_{t-1},s_{t-1},\cdots,s_{1})=\mathbb{P}(A_{t}=a_{t}|S_{t}=s_{t},A_{t-1}=a_{t-1},S_{t-1}=s_{t-1},\cdots,S_{1}=s_{1})$ in (4). Specifically, we require that for any two state sequences $\{s_{\ell}^{(1)}\}_{\ell}$ ,$\{s_{\ell}^{(2)}\}_{\ell}$ such that $\phi(s_{\ell}^{(1)})=\phi(s_{\ell}^{(2)})$, any action sequence $\{a_{\ell}\}_{\ell}$ and any time index $t\geq 1$,
@@
The following theorem validates the abstraction produced by the proposed DSA at any iteration.
Theorem 2 (OPE under the iterative procedure)
Finally, we note that the initialization of the iterative procedure doesn’t necessarily have to begin with forward abstraction; backward abstraction can serve as the starting point as well. In our experiments, both starting points have their merits, with their effectiveness varying depending on the environment. However, the overall differences in results are small.
@@ 5 Numerical experiments
Comparisons. We compare the proposed deep state abstraction (denoted by ‘DSA’) against single-iteration forward (‘forward’), backward (‘backward’) abstractions, the Markov state abstraction (Allen et al.,, 2021) (‘MSA’) and a reconstruction-based abstraction (Lange and Riedmiller,, 2010) (‘auto-encoder’). For fairer comparison, each abstraction’s performance is tested by applying a base FQE algorithm (Le et al.,, 2019) applied to the abstract state space. We also report the performance of FQE applied to the unabstracted, ground state space (‘FQE’).
Environment. We consider the “LunarLander-v2” environment in this section, with an original state dimension of 8. We manually include 292 irrelevant variables in the state, leading to a challenging 300-dimensional system. Refer to Appendix C for more details about the environment.
Results. We report MSEs and absolute biases of different post-abstraction-OPE estimators and those of the baseline FQE estimator without abstraction in Figure 5. It can be seen that the proposed DSA method outperforms other baseline methods, with the smallest relative MSE and absolute bias in most cases. To conclude, our analysis answers the following questions:
@@
    Is state abstraction useful for OPE? Both figures show that the baseline FQE applied to the ground state space performs the worst among all cases. This comparison reveals the usefulness of state abstractions for OPE.
@@
    Is the deep/iterative procedure more effective compared to single-iteration procedure? Notice that ‘Markov’ and ‘auto-encoder’ are types of model-irrelevant abstractions. The comparisons against these abstractions as well as ‘forward’ and ‘backward’ demonstrate the advantages of DSA over single-iteration procedures.
Figure 5: Relative MSEs and absolute biases of FQE estimators when applied to ground and abstract state spaces with various abstractions. The behavior policy is $\epsilon$-greedy with respect to the target policy, with $\epsilon=0.1,0.3,0.5$ from left to right.
@@
This appendix is structured as follows: Section A introduces additional related works on confounder selection in causal inference. The implementation details of the proposed state abstraction are discussed in Section B. Additional information concerning the environment and computing resources utilized is presented in Section C. All technical proofs can be found in Section D.
@@ Appendix A Confounder selection in causal inference
Broadly speaking, confounding refers to the problem that even if two variables are not causes of each other, they may exhibit statistical association due to common causes. Controlling for confounding is a central problem in the design of observational studies, and many criteria for confounder selection have been proposed in the literature. A commonly adopted criterion is the “common cause heuristic”, where the user only controls for covariates that are related to both the treatment and the outcome (Glymour et al.,, 2008; Austin,, 2011; Shortreed and Ertefaie,, 2017; Koch et al.,, 2020). Another widely used criterion is to simply use all covariates that are observed before the treatment in time (Rubin,, 2009; Hernán and Robins,, 2010, 2016). However, both of these approaches are not guaranteed to find a set of covariates that are sufficient to control for confounding. From a graphical perspective, confounder selection is essentially about finding a set of covariates that block all “back-door” paths (Pearl,, 2009), but this requires full structural knowledge about the causal relationship between the variables which is often not possible. This motivated some methods that only require partial structural knowledge (Vander Weele and Shpitser,, 2011; VanderWeele,, 2019; Guo and Zhao,, 2023). All the aforementioned methods need substantive knowledge about the treatment, outcome, and covariates. Other methods use statistical tests (usually of conditional independence) to trim a set of covariates that are assumed to control for confounding (Robins,, 1997; Greenland et al.,, 1999; Hernán and Robins,, 2010; De Luna et al.,, 2011; Belloni et al.,, 2014; Persson et al.,, 2017). The reader is referred to Guo et al., (2022) for a recent survey of objectives and approaches for confounder selection.

### v3

@@ 4.2 Iterative state abstraction
To elaborate, we assume for now that the behavior policy $b$ is not history-dependent. We will relax this condition later. Under the MDP model assumption, actions and states following $S_{t}$ are independent of those occurred prior to the realization of $S_{t}$. Accordingly, $(S_{t-1},A_{t-1})$ is conditionally independent of $\{(S_{k},A_{k})\}_{k>t}$ given $S_{t}$. Recall that $T$ corresponds to the termination time of trajectories in the offline dataset. We define a time-reversed MDP consisting of state-action-reward triplets $\{(S_{t},A_{t},b(A_{t}|S_{t})):t=T,\dots,1\}$. Its dynamics is described as follows (see also Figure 2(b) for the configuration):
@@
Given this MDP, the second type of MSA is readily available by applying the model-irrelevance condition to the reversed process. Analogous to Definition 5, our objective is to identify a state abstraction that is crucial for predicting the reward (e.g., the behavior policy) and the reversed transition function. We refer to the resulting condition as backward-model-irrelevance (see the formal definition below), since it is derived from the backward MDP model.
@@
The conditions of backward-model-irrelevance are similar to those specified for model-irrelevance outlined in Definition 5. Equation (6) essentially requires behavior-policy-irrelevance, or reward-irrelevance in the backward MDP, whereas Equation (7) corresponds to the “backward-transition-irrelevance”, and is equivalent to the conditional independence assumption between the pair $(A_{t},\phi(S_{t}))$ and $S_{t+1}$ given $\phi(S_{t+1})$.
@@
Next, we draw a connection between the abstraction derived from the backward-model-irrelevant condition and that developed by Allen et al. (2021) for policy learning. The latter imposes two conditions: (i) inverse-model-irrelevance, which requires $A_{t}$ to be conditionally independent of $S_{t}$ and $S_{t+1}$ given $\phi(S_{t})$ and $\phi(S_{t+1})$; (ii) density-ratio-irrelevance, which requires $\phi(S_{t})$ to be conditionally independent of $S_{t+1}$ given $\phi(S_{t+1})$. For effective policy learning, Allen et al. (2021) requires both conditions to hold in data generating processes following a diverse range of behavior policies. When restricting them to one behavior policy, (i) and (ii) imply the backward-transition-irrelevance condition in (7) whereas backward-transition-irrelevance in turn yields (ii). This allows us to adapt their algorithm to train our type-II MSA that satisfies backward-model-irrelevance; see Appendix B.2 for details.
@@
As mentioned earlier, due to the iterative nature of our procedure, after type-I MSA, the behavior policy when restricted to the abstract space can be history-dependent. To address this challenge, we modify the proposed backward-model-irrelevance by employing a history-dependent behavior policy $b_{t}(a_{t}|s_{t},a_{t-1},s_{t-1},\cdots,s_{1})=\mathbb{P}(A_{t}=a_{t}|S_{t}=s_{t},A_{t-1}=a_{t-1},S_{t-1}=s_{t-1},\cdots,S_{1}=s_{1})$ in (6). Specifically, we require that for any two state sequences $\{s_{\ell}^{(1)}\}_{\ell}$ ,$\{s_{\ell}^{(2)}\}_{\ell}$ such that $\phi(s_{\ell}^{(1)})=\phi(s_{\ell}^{(2)})$, any action sequence $\{a_{\ell}\}_{\ell}$ and any time index $t\geq 1$,
@@
The following lemma guarantees that the resulting abstraction $\phi$, obtained by applying the refined backward-model-irrelevance condition to the original MDP $(S_{t},A_{t},R_{t})_{t\geq 1}$ with a history-dependent behavior policy, remains an MSA.
Lemma 2
The abstraction $\phi$ derived from (7) and (8) is an MSA, regardless of whether the behavior policy is history-dependent or not.
Finally, given an abstraction derived from (7) and (8), we similarly augment it with $\{\pi(a|\bullet):a\in\mathcal{A}\}$ to ensure $\pi$-irrelevance.
Summary. To summarize, we have presented the two types of MSAs in this section. By iteratively alternating between these abstractions, the state space becomes progressively finer. As a result of our theory, the MSE of the post-abstraction OPE estimators is expected to decrease significantly. To conclude this section, we note that the initialization of the iterative procedure doesn’t necessarily have to begin with type-I MSA; type-II MSA can serve as the starting point as well. In our experiments, both starting points have their merits, with their effectiveness varying depending on the environment. However, the overall differences in results are small.
@@ 5 Numerical experiments
Comparisons. We compare the proposed iterative procedure, which yields a deeply abstracted state space (denoted by ‘DSA’, short for deep state abstraction), against the following baselines: single-iteration type-I MSA (‘Type-I’), type-II MSA (‘Type-II’), the abstract reward process (Chaudhari et al., 2024, ‘STAR’), the MSA proposed by Allen et al. (2021, ‘MSA’), and a reconstruction-based abstraction (Lange and Riedmiller, 2010, ‘Auto-Encoder’). For a fair comparison, except for STAR, which uses a built-in IS estimator for OPE, the performance of other abstractions is evaluated by applying a base FQE algorithm (Le et al., 2019) to the abstract state space. We also report the performance of FQE applied to the original ground state space (‘FQE’). This yields a total of six baselines, where:
  • Type-I and Type-II are variants of the proposed procedure;
  • FQE is a popular OPE method without any abstraction;
  • STAR is the state-of-the-art abstraction designed for OPE;
  • MSA and Auto-Encoder are state-of-the-art abstractions for policy learning.
The behavior policy is $\epsilon$-greedy with respect to the target policy, with $\epsilon=0.1,0.3,0.5$. The simulation results presented are averaged over 30 trials.
Environments. We consider four environments, including two complex games — “Atlantis-ram-v5" and “SpaceInvaders-ram-v5" — a medium-to-high difficulty environment, “LunarLander-v2", and a simple control problem, “Cartpole-v1". For “LunarLander-v2" and “Cartpole-v1", we manually introduce 292 and 296 irrelevant variables into the state, respectively, resulting in a challenging 300-dimensional system. Additional details about these environments can be found in Appendix C.
Results. We report the root MSE (RMSE) and root absolute bias of different post-abstraction OPE estimators, as well as the baseline FQE estimator without abstraction, in Figure 3. The results demonstrate that the proposed DSA generally outperforms other baseline methods, achieving the smallest RMSE in most cases across "Atlantis-ram-v5", "LunarLander-v2", and "Cartpole-v1". For "SpaceInvaders-ram-v5", DSA achieves the second smallest RMSE, while STAR performs the best in this environment. However, we note that STAR’s performance is highly unstable, as it performs the worst in the other three environments.
To conclude, our analysis answers the following questions:
@@
    Is state abstraction useful for OPE? Figure 3 shows that the baseline FQE applied to the ground state space performs worse than all post-abstraction FQE estimators. This comparison highlights the utility of state abstractions for OPE.
@@
    Is the iterative procedure more effective than single-iteration procedures? Comparisons against single-iteration abstractions (e.g., ‘Type-I’ and ‘Type-II’) demonstrate the advantages of multi-iteration procedures. This formally validates our theoretical finding that finer Markov state abstractions lead to lower MSE in the resulting OPE estimators.
  3. 3.
    Is the proposed abstraction better than other abstractions for OPE? In general, yes. The proposed DSA consistently achieves the best or second best performance across environments.
Figure 3: RMSEs and root absolute biases of various estimators for 4 different environments.
6 Conclusion
This paper develops an iterative procedure for state abstraction in OPE. Our theory suggests that finer Markov state abstractions generally lead to more accurate OPE estimators. Building on this insight, we propose an iterative procedure to maximally compress the state space while preserving the Markov property. Empirical investigations of our proposal’s performance confirm our theoretical findings.
There are certain limitations to our proposal: First, our theoretical analysis does not account for the estimation error incurred during the construction of the abstractions. Second, determining the optimal number of iterations, $K$, remains an open question. This choice involves a trade-off: a larger $K$ yields a smaller abstracted state space, but the estimation error for learning the abstraction may accumulate over iterations. In complex environments such as Atlantis-ram-v5 and SpaceInvaders-ram-v5, we set $K=4$ and observe good empirical performance. However, it remains unclear how to data-adaptively determine this hyperparameter.
@@ 6 Conclusion
This appendix is structured as follows: Section A introduces additional related works on confounder selection in causal inference. The implementation details of the proposed state abstraction are discussed in Section B. Additional information concerning the environment and computing resources utilized is presented in Section C. All technical proofs can be found in Section D. Toy examples demonstrating the advantages of DSA are presented in Section E, while additional conditions and properties of the proposed MSAs are discussed in Section F.
@@ Appendix A Confounder selection in causal inference
Broadly speaking, confounding refers to the problem that even if two variables are not causes of each other, they may exhibit statistical association due to common causes. Controlling for confounding is a central problem in the design of observational studies, and many criteria for confounder selection have been proposed in the literature. A commonly adopted criterion is the “common cause heuristic”, where the user only controls for covariates that are related to both the treatment and the outcome [Glymour et al., 2008, Austin, 2011, Shortreed and Ertefaie, 2017, Koch et al., 2020]. Another widely used criterion is to simply use all covariates that are observed before the treatment in time [Rubin, 2009, Hernán and Robins, 2010, 2016]. However, both of these approaches are not guaranteed to find a set of covariates that are sufficient to control for confounding. From a graphical perspective, confounder selection is essentially about finding a set of covariates that block all “back-door” paths [Pearl, 2009], but this requires full structural knowledge about the causal relationship between the variables which is often not possible. This motivated some methods that only require partial structural knowledge [Vander Weele and Shpitser, 2011, VanderWeele, 2019, Guo and Zhao, 2023]. All the aforementioned methods need substantive knowledge about the treatment, outcome, and covariates. Other methods use statistical tests (usually of conditional independence) to trim a set of covariates that are assumed to control for confounding [Robins, 1997, Greenland et al., 1999, Hernán and Robins, 2010, De Luna et al., 2011, Belloni et al., 2014, Persson et al., 2017]. The reader is referred to Guo et al. [2022] for a recent survey of objectives and approaches for confounder selection.